# Methodology for constraining ultralight vector bosons with gravitational wave searches targeting merger remnant black holes

### v1

@@ I Introduction
Through the superradiance mechanism, ultralight bosons can form gravitationally bound field configurations around a spinning BH, growing to form a macroscopic cloud as mass-energy is extracted from the BH Penrose (1969); Press and Teukolsky (1972); Zel’Dovich (1971); Starobinskii (1973); Detweiler (1980); Bekenstein (1973); Dolan (2007); Arvanitaki and Dubovsky (2011). The superradiant mechanism of a bosonic field configuration of frequency $\omega$ and azimuthal number $m$ operates efficiently when two conditions are met: a) The BH’s horizon frequency $\Omega_{H}$ is greater than that of the bosonic field configuration, $\omega<m\Omega_{H}$, and b) the Compton wavelength of the boson is comparable to the BH’s horizon radius. The latter implies that the dimensionless coupling constant $\alpha=GMm_{b}$, which links the BH mass $M$ and boson mass $m_{b}$, satisfies $\alpha\lesssim\mathcal{O}(1)$. (Here and in the following, we employ $c=1$ and $\hbar=1$ units.) Systems meeting these criteria undergo a phase of exponential growth of the boson cloud, saturating gravitationally by spinning down the BH until condition a) is roughly saturated. What remains is a boson cloud with a mass of at most 10% of the BH’s mass, which dissipates as it emits gravitational waves (GWs) Arvanitaki and Dubovsky (2011); Yoshino and Kodama (2014, 2015); Arvanitaki et al. (2015, 2017); Brito et al. (2017a, b); Yoshino and Kodama (2014, 2015); Baryakhtar et al. (2017); Chan and Hannuksela (2024); Cardoso et al. (2018); Hannuksela et al. (2019); Zhang and Yang (2019); East (2017); East and Pretorius (2017); East (2018); Siemonsen and East (2020); Brito et al. (2015). This radiation is quasi-monochromatic with a frequency of roughly twice the oscillation frequency of the boson cloud Arvanitaki et al. (2015) and a slight upward frequency drift Baryakhtar et al. (2017). While the superradiant instability is in principle possible for a BH of any mass, granted a boson particle of appropriate mass exists, our interest lies with stellar-mass BHs, whose matching bosons emit at frequencies falling within the sensitive band of both existing ground-based GW detectors (i.e., the Advanced Laser Interferometer Gravitational-Wave Observatory (aLIGO) Aasi et al. (2015), Advanced Virgo Acernese et al. (2014), and KAGRA Akutsu et al. (2021)) and future ones (e.g., Einstein Telescope Punturo et al. (2010); Hild et al. (2011); Team (2011); Maggiore et al. (2020) and Cosmic Explorer Abbott et al. (2017a); Evans et al. (2021); Reitze et al. (2019)). In targeting stellar-mass BHs, we can probe a boson mass range from roughly $10^{-14}$ eV to $10^{-11}$ eV Arvanitaki et al. (2015); Brito et al. (2017b).
@@
Reference Isi et al. (2019) shows that we can use continuous wave (CW) analysis techniques to search for signals from scalar clouds around merger remnant BHs. However, these targets generally lie outside the sensitivity range of current-generation detectors. In Ref. Jones et al. (2023), it is demonstrated that vector clouds around merger remnant BHs, which emit stronger but shorter GW signals compared to scalar clouds, can potentially be reached with current-generation detectors using techniques adapted from conventional CW searches. The directed search method in Ref. Jones et al. (2023) for targeting vector signals from merger remnant BHs is adapted from the semicoherent hidden Markov model (HMM) search method used in more traditional CW searches (see, e.g., Refs. Suvorova et al. (2016); Sun et al. (2018); Isi et al. (2019); Sun and Melatos (2019)). Reference Jones et al. (2023) also assumes that the parameters of the target BH (the mass, spin, sky position, and orientation) are perfectly known, so that for a given boson mass, one can determine the expected GW signal with a sufficiently good model Siemonsen et al. (2023a). In this case, constraining the boson mass in the absence of a detection is fairly straightforward. However, in reality, no system parameter is perfectly known; the best knowledge of the remnant BH comes from the posterior distribution obtained through compact object coalescence (CBC) analysis of the merger event Abbott et al. (2016a); Veitch et al. (2015). In this paper, we formulate a procedure to incorporate these parameter uncertainties when inferring constraints on the boson mass in the absence of a detection.
@@ II.1 Long-transient search method
The search method Jones et al. (2023) is summarized as follows. We take the short Fourier transforms (SFTs) of the time series data from a given GW detector over the relevant stretch of observing time $T_{\rm obs}$ during which we expect a signal may be detectable. We use these SFTs to compute a frequency-domain matched filter ($\mathcal{F}$-statistic) over a coherent time interval $T_{\rm coh}$, set by the frequency drift of the signal. Then, we combine these segments of length $T_{\rm coh}$ incoherently with the HMM-based tracking algorithm. We define the detection statistic as the total log-likelihood of the optimal path returned by the HMM algorithm, divided by the number of coherent segments in the search, i.e., Eq. (24) in Ref. Jones et al. (2023). Candidates with a detection statistic above the threshold for a chosen false alarm probability, $P_{\rm fa}$, are identified and followed up using some kind of validation process similar to what has been done in previous CW searches (see, e.g., Refs. Abbott et al. (2022c, 2021b)).
@@ II.2 CBC parameter estimation
We start by briefly summarizing how the BH parameter uncertainties we must account for arise. When one or more detectors identify a candidate signal, different types of searches are run to determine if it is a real GW signal or simply noise. One type of search attempts to recover the signal assuming the candidate was the result of a CBC Abbott et al. (2016b). It uses optimal matched filtering with a template bank of waveforms predicted by general relativity. The other type of search does not make any particular assumptions about the signal waveform; instead, it considers a variety of generic transient signals Klimenko et al. (2008, 2016). Although the methods used in these two types of searches are independent, we would expect sufficiently loud CBC signals to be recovered by both Abbott et al. (2016a).
@@ II.3 Constraining the boson mass
In a typical CW search, we usually derive the upper limits on the signal strain (and the corresponding source properties) in the absence of a detection in a frequentist way, as outlined in, e.g., Ref. Abbott et al. (2022c), and summarized below. We start by considering a sub-band frequency range (e.g., 1 Hz) within which the search is carried out, and we assume no signal has been detected. The target source is characterized by a multi-dimensional vector parameter $\bm{\theta}=\{{\theta}_{1},{\theta}_{2},\cdots,{\theta}_{M}\}$, where each component parameter is associated with a distribution. Then, we generate $N$ synthetic signals, each based on a randomly drawn set of system parameters $\bm{\theta}_{i}=\{{\theta_{1}}_{i},{\theta_{2}}_{i},\cdots,{\theta_{M}}_{i}\}$ with $i=1,2,\cdots,N$ and inject them into $N$ noise realizations. We repeat such injection study by setting the strain $h_{0}$ to different values in order to find the strain value $h_{0}^{\rm CL}$, with the superscript $\rm CL$ indicating the ratio of the injections recovered (for example, when ${\rm CL}=95\%$, the value of $h_{0}^{95\%}$ indicates the “95% confidence upper limit” derived in a given sub-band).
@@
where $N_{\rm det}(\bm{\theta}_{i};m_{V})$ is the number of recovered signals in $N_{\rm noise}$ noise realizations for a set of given BH parameters and can take on values between $[0,N_{\rm noise}]$. In the absence of a detection in the search, $P_{\rm det}(m_{V})$ corresponds to the confidence level with which we can exclude the existence of vectors with mass $m_{V}$. We describe this procedure through detailed simulations in the following section.
@@ III.1 Number of BH samples
In a typical set of parameter estimation results of a CBC event, e.g., from the LIGO-Virgo-KAGRA catalogs Abbott et al. (2019, 2021a, 2023), there are $\mathcal{O}(10^{4})$ posterior samples which form the full multi-dimensional posterior distribution. If we take all the posterior samples of the remnant BH, we have $N_{\rm BH}\sim 10^{4}$ in Eq. (5). For each set of sampled system parameters $\bm{\theta}_{i}$, corresponding to a particular definitive BH target, we usually quantify the recovery rate by simulating the system and injecting the signal into multiple noise realizations (e.g., $N_{\rm noise}\sim 10^{2}$), as described in Sec. II.3. To derive the constraints on $m_{V}$, we need to repeat the simulation procedure, calculating $P_{\rm det}(m_{V})$ over a range of $m_{V}$ values in the most accessible parameter space, corresponding to $\sim 10$ values of $m_{V}$. Thus, in the worst case, $P_{\rm det}$ needs to be computed using a frequentist approach with $\mathcal{O}(10^{7})$ trials to include both BH-boson system configurations and noise realizations, which is computationally infeasible. Instead, we randomly sample $N$ sets of parameters from the BH posterior distribution and quantify exactly how large $N$ needs to be to such that the subset of samples represents the full distribution reasonably well.
@@ III.2 Synthetic system and constraints
For a given CBC remnant BH, let us first determine a range of $m_{V}$ to be studied. Even with a single set of BH parameters, a range of boson masses can be tested. This is because bosons with masses close to the value that optimally matches the sample BH can yield GW emission at a nearly maximized strength and have equally good observational prospects (e.g., see Fig. 11 in Ref. Jones et al. (2023)). However, as discussed above, the BH parameters also have associated uncertainties which we must consider. In general, we can take $m_{V}\approx[0.6,1.1]m_{V}^{\rm opt}$ as the promising boson mass range, where $m_{V}^{\rm opt}$ here is the boson mass which maximizes the strain amplitude for the median values of the BH parameters. In principle, for each sampled set of BH parameters, a different optimally matched boson mass can be obtained. For BHs with masses within two standard deviations of the mean of the mass distribution, all corresponding optimally matched bosons should fall within the above vector mass range. For the synthetic system (with median parameters $M=53.2\ M_{\odot}$ and $\chi=0.72$), we have $m_{V}^{\rm opt}=4.612\times 10^{-13}$ eV, so the promising search range is $m_{V}=[2.767,5.073]\times 10^{-13}$ eV. We take a sample of 11 evenly-spaced values from across this range, $m_{V}=\{2.767,2.998,3.228,3.459,3.689,3.920,4.151,4.381,4.612,\\ 4.842,5.073\}\times 10^{-13}$ eV, and we calculate the detection probability $P_{\rm det}(m_{V})$ for each $m_{Vj}$ (where the subscript $j$ indexes the $m_{V}$ value in the array) according to the guidelines that follow.
@@
To avoid bias caused by any assumptions made about the target BH’s true parameters, for a given $m_{V}$ value, we randomly draw $N_{\rm BH}=200$ sample parameter sets from the distribution of the GW170814-like remnant BH (Table 1). We generate a synthetic signal based on each sample set of system parameters $\bm{\theta}_{i}$; $m_{Vj}$}, and then we inject each signal into $N_{\rm noise}=10$ independent Gaussian noise realizations. This injection procedure is repeated for each of the $N_{\rm BH}=200$ samples with parameters $\bm{\theta}_{i}$. When attempting to recover each injection, we must use the same search configurations that would have been used in the search run on real detector data. In practice, this looks like a grid of configuration setups (search start time $t_{\rm start}$, coherent time $T_{\rm coh}$, and total observing time $T_{\rm obs}$) chosen based on the interesting parameter space. For details on how we choose the configurations to adequately cover the parameter space, see Appendix A. In this example, we have 11 configurations. For any synthetic signal injection, it is considered a successful recovery as long as the injection is recovered by at least one configuration. Using all 11 configurations, we calculate the recovery rate $P_{\rm det}(m_{V_{j}})$ for a given $m_{V_{j}}$, out of $N_{\rm BH}N_{\rm noise}=2000$ injections [Eq. (5)]. [Note, due to the precision of $P_{\rm det}(m_{V_{j}})$, further increasing either the number of samples $N_{\rm BH}$ or the number of noise realizations $N_{\rm noise}$ is not beneficial.] We repeat this process for each value of $m_{V_{j}}$.
@@
Figure 2: The detection probability $P_{\rm det}$ as a function of $m_{V}$ with a 1% (orange), 5% (blue), and 10% (purple) false alarm probability. The vertical dashed line marks the optimal boson mass $m_{V}^{\rm opt}$ for the GW170814-like synthetic system with parameters shown in the median row in Table 1. The dark gray (light gray) shaded region indicates the disfavored boson mass range with 80% (50%) confidence ($P_{\rm fa}=1\%$), assuming a search targeting such a remnant BH does not yield a detection.

### v2

@@ I Introduction
Through the superradiance mechanism, ultralight bosons can form gravitationally bound field configurations around a spinning BH, growing to form a macroscopic cloud as mass-energy is extracted from the BH Penrose (1969); Press and Teukolsky (1972); Zel’Dovich (1971); Starobinskii (1973); Detweiler (1980); Bekenstein (1973); Dolan (2007); Arvanitaki and Dubovsky (2011). The superradiant mechanism of a bosonic field configuration of frequency $\omega$ and azimuthal number $m$ operates efficiently when two conditions are met: a) The BH’s horizon frequency $\Omega_{H}$ is greater than that of the bosonic field configuration, $\omega<m\Omega_{H}$, and b) the Compton wavelength of the boson is comparable to the BH’s horizon radius. The latter implies that the dimensionless coupling constant $\alpha=GMm_{b}$, which links the BH mass $M$ and boson mass $m_{b}$, satisfies $\alpha\lesssim\mathcal{O}(1)$. (Here and in the following, we employ $c=1$ and $\hbar=1$ units.) Systems meeting these criteria undergo a phase of exponential growth of the boson cloud, saturating gravitationally by spinning down the BH until condition a) is roughly saturated. What remains is a boson cloud with a mass of at most 10% of the BH’s mass, which dissipates as it emits gravitational waves (GWs) Arvanitaki and Dubovsky (2011); Yoshino and Kodama (2014, 2015); Arvanitaki et al. (2015, 2017); Brito et al. (2017a, b); Yoshino and Kodama (2014, 2015); Baryakhtar et al. (2017); Chan and Hannuksela (2024); Cardoso et al. (2018); Hannuksela et al. (2019); Zhang and Yang (2019); East (2017); East and Pretorius (2017); East (2018); Siemonsen and East (2020); Brito et al. (2015). This radiation is quasi-monochromatic with a frequency of roughly twice the oscillation frequency of the boson cloud Arvanitaki et al. (2015) and a slight upward frequency drift Baryakhtar et al. (2017). While the superradiant instability is in principle possible for a BH of any mass, granted a boson particle of appropriate mass exists, our interest lies with stellar-mass BHs, whose matching bosons emit at frequencies falling within the sensitive band of both existing ground-based GW detectors (i.e., the Advanced Laser Interferometer Gravitational-Wave Observatory (aLIGO) Aasi et al. (2015), Advanced Virgo Acernese et al. (2014), and KAGRA Akutsu et al. (2021)) and future ones (e.g., Einstein Telescope Punturo et al. (2010); Hild et al. (2011); Team (2011); Maggiore et al. (2020) and Cosmic Explorer Abbott et al. (2017a); Evans et al. (2021); Reitze et al. (2019)). In targeting stellar-mass BHs, one can probe a boson mass range from roughly $10^{-14}$ eV to $10^{-11}$ eV Arvanitaki et al. (2015); Brito et al. (2017b).
@@
Reference Isi et al. (2019) shows that one can use continuous wave (CW) analysis techniques, which were initially designed to search for gravitational waves from spinning non-axisymmetric neutron stars, to search for signals from scalar clouds around merger remnant BHs. However, these targets generally lie outside the sensitivity range of current-generation detectors. In Ref. Jones et al. (2023), it is demonstrated that vector clouds around merger remnant BHs, which emit stronger but shorter GW signals compared to scalar clouds, can potentially be reached with current-generation detectors using techniques adapted from conventional CW searches. The directed search method in Ref. Jones et al. (2023) for targeting vector signals from merger remnant BHs is adapted from the semicoherent hidden Markov model (HMM) search method used in more traditional CW searches (see, e.g., Refs. Suvorova et al. (2016); Sun et al. (2018); Isi et al. (2019); Sun and Melatos (2019)). Reference Jones et al. (2023) also assumes that the parameters of the target BH (the mass, spin, sky position, and orientation) are perfectly known, so that for a given boson mass, one can determine the expected GW signal with a sufficiently good model Siemonsen et al. (2023a). In this case, constraining the boson mass in the absence of a detection is fairly straightforward. However, in reality, no system parameter is perfectly known; the best knowledge of the remnant BH comes from the posterior distribution obtained through compact object coalescence (CBC) analysis of the merger event Abbott et al. (2016a); Veitch et al. (2015). In this paper, we formulate a procedure to incorporate these parameter uncertainties when inferring constraints on the boson mass in the absence of a detection.
@@ II.1 Long-transient search method
The search method Jones et al. (2023) is summarized as follows. We take the short Fourier transforms (SFTs) of the time series data from a given GW detector over the relevant stretch of observing time $T_{\rm obs}$ during which we expect a signal may be detectable. We use these SFTs to compute a frequency-domain matched filter ($\mathcal{F}$-statistic) over a coherent time interval $T_{\rm coh}$, set by the frequency drift of the signal. Then, we combine these segments of length $T_{\rm coh}$ incoherently with the HMM-based tracking algorithm. We define the detection statistic as the total log-likelihood of the optimal path returned by the HMM algorithm, divided by the number of coherent segments in the search, i.e., Eq. (24) in Ref. Jones et al. (2023). Candidates with a detection statistic above the threshold for a chosen false alarm probability, $P_{\rm fa}$, are identified and followed up using a similar kind of validation process to what has been done in previous CW searches (see, e.g., Refs. Abbott et al. (2022c, 2021b)).
@@ II.2 CBC parameter estimation
We start by briefly summarizing how the BH parameter uncertainties that must be accounted for arise. When one or more detectors identify a candidate signal, different types of searches are run to determine if it is a real GW signal or simply noise. One type of search attempts to recover the signal assuming the candidate was the result of a CBC Abbott et al. (2016b). It uses optimal matched filtering with a template bank of waveforms predicted by general relativity. The other type of search does not make any particular assumptions about the signal waveform; instead, it considers a variety of generic transient signals Klimenko et al. (2008, 2016). Although the methods used in these two types of searches are independent, we would expect sufficiently loud CBC signals to be recovered by both Abbott et al. (2016a).
@@ II.3 Constraining the boson mass
In a typical CW search, the upper limits on the signal strain (and the corresponding source properties) in the absence of a detection are usually derived in a frequentist way, as outlined in, e.g., Ref. Abbott et al. (2022c), and summarized below. We start by considering a sub-band frequency range (e.g., 1 Hz) within which the search is carried out, and we assume no signal has been detected. The target source is characterized by a multi-dimensional vector parameter $\bm{\theta}=\{{\theta}_{1},{\theta}_{2},\cdots,{\theta}_{M}\}$, where each component parameter is associated with a distribution. Then, we generate $N$ synthetic signals, each based on a randomly drawn set of system parameters $\bm{\theta}_{i}=\{{\theta_{1}}_{i},{\theta_{2}}_{i},\cdots,{\theta_{M}}_{i}\}$ with $i=1,2,\cdots,N$ and inject them into $N$ noise realizations. We repeat such injection study by setting the strain $h_{0}$ to different values in order to find the strain value $h_{0}^{\rm CL}$, with the superscript $\rm CL$ indicating the ratio of the injections recovered (for example, when ${\rm CL}=95\%$, the value of $h_{0}^{95\%}$ indicates the “95% confidence upper limit” derived in a given sub-band).
@@
where $N_{\rm det}(\bm{\theta}_{i};m_{V})$ is the number of recovered signals in $N_{\rm noise}$ noise realizations for a set of given BH parameters and can take on values between $[0,N_{\rm noise}]$. In the absence of a detection in the search, $P_{\rm det}(m_{V})$ corresponds to the confidence level with which one can exclude the existence of vectors with mass $m_{V}$. We describe this procedure through detailed simulations in the following section.
@@ III.1 Number of BH samples
In a typical set of parameter estimation results of a CBC event, e.g., from the LIGO-Virgo-KAGRA catalogs Abbott et al. (2019, 2021a, 2023), there are $\mathcal{O}(10^{4})$ posterior samples which form the full multi-dimensional posterior distribution. If we take all the posterior samples of the remnant BH, this yields $N_{\rm BH}\sim 10^{4}$ in Eq. (5). For each set of sampled system parameters $\bm{\theta}_{i}$, corresponding to a particular definitive BH target, we usually quantify the recovery rate by simulating the system and injecting the signal into multiple noise realizations (e.g., $N_{\rm noise}\sim 10^{2}$), as described in Sec. II.3. To derive the constraints on $m_{V}$, the simulation procedure should be repeated, calculating $P_{\rm det}(m_{V})$ over a range of $m_{V}$ values in the most accessible parameter space, corresponding to $\sim 10$ values of $m_{V}$. Thus, in the worst case, $P_{\rm det}$ needs to be computed using a frequentist approach with $\mathcal{O}(10^{7})$ trials to include both BH-boson system configurations and noise realizations, which is computationally infeasible. Instead, we randomly sample $N$ sets of parameters from the BH posterior distribution and quantify exactly how large $N$ needs to be to such that the subset of samples represents the full distribution reasonably well.
@@ III.2 Synthetic system and constraints
For a given CBC remnant BH, let us first determine a range of $m_{V}$ to be studied. Even with a single set of BH parameters, a range of boson masses can be tested. This is because bosons with masses close to the value that optimally matches the sample BH can yield GW emission at a nearly maximized strength and have equally good observational prospects (e.g., see Fig. 11 in Ref. Jones et al. (2023)). However, as discussed above, the BH parameters also have associated uncertainties which must be considered. In general, one can take $m_{V}\approx[0.6,1.1]m_{V}^{\rm opt}$ as the promising boson mass range, where $m_{V}^{\rm opt}$ here is the boson mass which maximizes the strain amplitude for the median values of the BH parameters. In principle, for each sampled set of BH parameters, a different optimally matched boson mass can be obtained. The above vector mass range has been chosen empirically based on the estimated signal strength produced by bosons with different masses, which is also validated by Fig. 11 of Ref. Jones et al. (2023), in which the bright region corresponds to the boson mass range that produces potentially detectable signals for a given BH. For BHs with masses within two standard deviations of the mean of the mass distribution, all corresponding optimally matched bosons should fall within this range. For the synthetic system (with median parameters $M=53.2\ M_{\odot}$ and $\chi=0.72$), we have $m_{V}^{\rm opt}=4.612\times 10^{-13}$ eV, so the promising search range is $m_{V}=[2.767,5.073]\times 10^{-13}$ eV. We take a sample of 11 evenly-spaced values from across this range, $m_{V}=\{2.767,2.998,3.228,3.459,3.689,3.920,4.151,4.381,4.612,\\ 4.842,5.073\}\times 10^{-13}$ eV, and we calculate the detection probability $P_{\rm det}(m_{V})$ for each $m_{Vj}$ (where the subscript $j$ indexes the $m_{V}$ value in the array) according to the guidelines that follow.
@@
To avoid bias caused by any assumptions made about the target BH’s true parameters, for a given $m_{V}$ value, we randomly draw $N_{\rm BH}=200$ sample parameter sets from the distribution of the GW170814-like remnant BH (Table 1). We generate a synthetic signal based on each sample set of system parameters $\bm{\theta}_{i}$; $m_{Vj}$}, and then we inject each signal into $N_{\rm noise}=10$ independent Gaussian noise realizations using the simulateCW Python module in the LALPulsar library of LALSuite LIGO Scientific Collaboration and Virgo Collaboration and KAGRA Collaboration (2018); Wette (2020). This injection procedure is repeated for each of the $N_{\rm BH}=200$ samples with parameters $\bm{\theta}_{i}$. When attempting to recover each injection, one must use the same search configurations that would have been used in the search run on real detector data. In practice, this looks like a grid of configuration setups (search start time $t_{\rm start}$, coherent time $T_{\rm coh}$, and total observing time $T_{\rm obs}$) chosen based on the interesting parameter space. For details on how to choose the configurations to adequately cover the parameter space, see Appendix A. In this example, we have 11 configurations. For any synthetic signal injection, it is considered a successful recovery as long as the injection is recovered by at least one configuration. Using all 11 configurations, we calculate the recovery rate $P_{\rm det}(m_{V_{j}})$ for a given $m_{V_{j}}$, out of $N_{\rm BH}N_{\rm noise}=2000$ injections [Eq. (5)]. [Note, due to the precision of $P_{\rm det}(m_{V_{j}})$, further increasing either the number of samples $N_{\rm BH}$ or the number of noise realizations $N_{\rm noise}$ is not beneficial.] We repeat this process for each value of $m_{V_{j}}$.
@@
One might notice that the curves in Fig. 2 all plateau across certain boson mass ranges. These plateaus are the result of a balancing act that occurs between the intrinsic strain amplitude of the signal, which decreases as the boson mass is shifted to lower values, and the search sensitivity, which improves for lower boson masses that produce signals with smaller frequency drift and thus allow larger $T_{\rm coh}$ values to be used (also described in Ref. Jones et al. (2023)). One can see a similar behavior in Fig. 11 of Ref. Jones et al. (2023), where the horizon distance peaks gradually over a range of vector masses rather than sharply over a single mass.
Figure 2: The detection probability $P_{\rm det}$ as a function of $m_{V}$ with a 1% (orange), 5% (blue), and 10% (purple) false alarm probability. The vertical dashed line marks the optimal boson mass $m_{V}^{\rm opt}$ for the GW170814-like synthetic system with parameters shown in the median row in Table 1. The dark gray (light gray) shaded region indicates the disfavored boson mass range with 80% (50%) confidence ($P_{\rm fa}=1\%$), assuming a search targeting such a remnant BH does not yield a detection. The error bars indicate the $1\sigma$ beta-binomial uncertainty in $P_{\rm det}$ Johnson et al. (2005).